# Prompting Large Language Models for Recommender Systems: A Comprehensive Framework and Empirical Analysis

### v1

@@ 5.2.4. Research Questions and Observations for Modeling Short-term Intentions
As shown in Table 6-7, we summarize the following findings of modeling short-term intentions for prompting LLMs.
@@
(1) The retrieval strategies for recent and relevant items (RQ1). First we conduct an experiment to explore the retrieval strategies for recent and relevant items, and evaluate the selection of query and memory for item retrieval. As shown in Table 6, we compare four combinations with queries and memories. In terms of the query, combining the design of long-term and short-term interest yields better results than only the short-term interest. When selecting items that represent interest for recommendation, the query needs to consider both recent tendencies and long-term preferences of users. Our results have shown that the personalized long-term interest of users can have a positive effect on retrieval for interest modeling. In terms of the memory, the design of the personalized memory with user interest outperforms the global memory with general descriptions. This result is in line with features of personalized recommender systems (Shu et al., 2023; Wang et al., 2023c), indicating the importance of personalization when designing the user interest for LLMs (Zhang et al., 2023c).
(2) The impact of different short-term intentions modeling forms (RQ2). To examine the effectiveness of various modeling forms w.r.t. the short-term interest, we design four variants based on the recently interacted 10 items and report their recommendation performance in Table 7. We can see that personalized memory can provide more information to LLMs than simple titles (comparison between the first and third lines), demonstrating the role of personalized memories with the user interest. On the other hand, we cannot completely replace the title with a description, as retaining the title and adding personalized descriptions perform even better results (comparison between the last two lines). Without the personalized memory, the least-to-most prompting strategy that allows LLMs to summarize recent items first can provide personalized descriptions and achieve performance improvement (comparison between the first two lines). There is also a difference in the impact of the short-term interest modeling forms on recommendation abilities between the two datasets, i.e., MovieLens and Amazon-Books. For the short-term intentions in the MovieLens dataset, the summarized text induced by LLMs is better than the personalized memory, while in the Amazon-Books dataset, results are opposite (comparison between the second and fourth lines). It is because there is no review comments but only the ratings in the MovieLens dataset, so capabilities of the personalized memory is relatively poor.
Observations of Modeling Short-term Intentions While selecting recent and relevant items, it is preferable to combine long-term and short-term user profiles with personalized descriptions to construct the personalized query and memory. For short-term interest of users, representing items with titles and generated personalized descriptions can further improve the effectiveness of LLMs for recommendation.
5.2.5. Research Questions and Observations for Modeling Long-term Preferences
In this section, we conduct experiments for modeling the long-term preferences of users in ProLLM4Rec. Specifically, we focus on the two research questions.
  • RQ1: Whether different modeling forms of long-term preferences have similar recommendation performance?
  • RQ2: How to combine short-term intentions and long-term preferences for the hybrid interest?
To evaluate the recommendation performance of different modeling methods for long-term preferences, we design the following variants of interest modeling forms to answer the first research question:
  • Recent items: it is the basic modeling form of selecting only recent 10 items for interest modeling.
  • Summarized user profile + recent items: it concatenates the user profile summarized by LLMs and the recently interacted 10 items for interest modeling, considering both the long-term and short-term interest.
  • Retrieved items from long-term history: it retrieves 10 relevant items from the long-term interest. Since recent items have a higher probability to be retrieved, this variant considers recency as well.
  • Retrieved items from long-term history + recent items: it uses both the retrieved 10 items from long-term history and the recent 10 items as the representation of user interest, considering both the long-term and short-term interest. It is worth noting that the retrieved items in this variant do not include the latest 10 items.
  • Retrieved personalized text from long-term history: it also retrieves 10 relevant items from the long-term history like “retrieved items from long-term history”, but this variant uses the personalized descriptions of items from the personalized memory instead of pure titles to represent the user interest.
  • Summarized text based on retrieved personalized text: it employs the retrieved personalized text from long-term history in the last variant by summarizing the retrieved 10 texts into one description. Not providing 10 specific items for LLMs, this variant only summarizes the user interest through personalized text.
Due to the positive impact of both long-term and short-term interest on the recommendation results of LLMs, we further explore different combination approaches of long-term and short-term interest to answer the second research question. Specifically, we consider the following four variants of the combination strategies.
  • Short-term interest only: it only utilizes the recently interacted 10 items as the baseline for comparison.
  • Concatenate long- and short-term interest: it combines the long-term and short-term interest by concatenating them together for user interest modeling.
  • Short-term as query for long-term interest: it combines the long-term and short-term interest by utilizing the recently interacted 10 items as the query for retrieving relevant items from the long-term history.
  • User profile and short-term as query for long-term interest: it retrieves recent and relevant items from the long-term interest, and for the query, both the personalized user profile and short-term interest are utilized.
Table 8. Performance comparison on various modeling forms of long-term interest.
User interest modeling MovieLens-1M Amazon-Books
ndcg@1 ndcg@10 ndcg@1 ndcg@10
recent items 0.1817 0.3985 0.2467 0.4276
summarized user profile + recent items 0.2000 0.4466 0.2500 0.4925
retrieved items from long-term history 0.1950 0.4343 0.2850 0.4650
retrieved items from long-term history + recent items 0.2400 0.4441 0.3250 0.4636
retrieved personalized text from long-term history 0.2550 0.4714 0.3150 0.5241
summarized text based on retrieved personalized text 0.2000 0.4492 0.2500 0.4872
Table 9. Performance comparison on the combination strategies of long- and short-term interest.
Combination strategies MovieLens-1M Amazon-Books
ndcg@1 ndcg@10 ndcg@1 ndcg@10
short-term interest only 0.1817 0.3985 0.2467 0.4276
concatenate long- and short-term interest 0.2000 0.4466 0.2500 0.4925
short-term as query for long-term interest 0.2400 0.4441 0.3250 0.4636
user profile and short-term as query for long-term interest 0.2400 0.4563 0.3300 0.4712
As shown in Table 8-9, key findings of the design for long-term preferences in ProLLM4Rec are summarized as follows.
(1) The impact of different long-term preferences modeling forms (RQ1). As for user interest modeling forms in the first research question, Table 8 illustrates the performance comparison on six modeling forms of the long-term interest. Through the comparison between the first two lines, we can see that long-term preferences of users can improve modeling effects of the short-term interest. Besides, modeling items retrieved from the long-term interest also improves the recommendation performance compared to only recent items (comparison between the first and third lines), and combining the retrieved items and recent item yields even better results in line four. On the whole, the first four lines in Table 8 indicate the key role of combining the long-term and short-term interest. Furthermore, retrieved personalized texts from the long-term memory greatly improve the recommendation performance compared to only using recent items. However, we further summarize the retrieved personalized text and results deteriorate, indicating that the rich information contained in the original sequence cannot be simply included as general descriptions. The last two lines in Table 8 highlight the important role of personalized memories w.r.t. personalized users.
(2) The combination strategies for the hybrid interest (RQ2). As for combination strategies of the long-term and short-term interest, we can see from Table 9 that simply concatenating the long-term and short-term sequences is the most commonly used and simple way, but not the optimal solution. One possible combination approach is to use the short-term interest as queries to retrieve the long-term interest, and empirical analysis shows better results. Performance comparison on the combination strategies validates the importance of combining the long-term and short-term interest.
Observations of Modeling Long-term Preferences For long-term interest, it is recommended to provide a personalized memory to store user preferences. Compared to the simple concatenation, retrieving long-term preferences based on short-term intentions is a better way for forming hybrid interest descriptions, and more approaches can be further explored.
@@ 5.3.1. Procedures of Candidate Items Construction
When using LLMs as a recommendation model, a crucial problem is that LLMs tend to recommend items not in the dataset (Geng et al., 2022). One of the approaches is to provide a limited number of candidate items for model selection. On the one hand, the source of candidate items determines the upper limit of personalized recommendation accuracy (Hou et al., 2023; Zhang et al., 2023f). On the other hand, the form of candidate items reflects the specific recommendation task and exerts huge impacts on performance of LLMs for recommendation (Dai et al., 2023a). As shown in Fig. 9, we will discuss the main procedures of processing candidate items, including the source, representation and grounding of candidate items, and then conduct experiments on the effects of candidate items in our framework ProLLM4Rec.
@@
Figure 9. Main procedures of candidate items construction in  ProLLM4Rec.
@@
$\bullet$ Source of Candidate Items. In practical industrial applications, recommendation is a multi-stage process, including two kinds of typical stages, i.e., the recall stage and the re-ranking stage (Bao et al., 2023a; Guo et al., 2017; Rendle et al., 2009). The recall stage in recommender systems is used for preliminary filtering and selection from the entire set of item candidates, while the re-ranking stage is used for readjusting the selected candidates, which is more suitable for LLMs as the re-ranker (Hou et al., 2023; Ma et al., 2023). Therefore, the source of candidate items to be ranked is crucial for the results of LLMs for recommendation. Popular approaches for selection of candidate items include traditional recommendation models (Yue et al., 2023) and retrieval algorithms (Lin et al., 2023b), both of which provide a list of potentially recommended items in a comparably coarse granularity.
@@
$\bullet$ Representation of Candidate Items. In the construction of prompts, we need to use historical items to represent users and provide candidate items for LLMs to rank, both of which involve item indexing (Geng et al., 2022; Lin et al., 2023c). However, there is a gap between the item representation in LLMs and recommendation, so the indexing of items between LLMs and recommender systems becomes the key and difficult point of ProLLM4Rec. There are three typical ways to index items: 1) token-based identifiers, 2) description-based identifiers and 3) hybrid identifiers (Zheng et al., 2023). For token-based identifiers, researchers often use numerical IDs to identify items (Geng et al., 2022). Since token-based identifiers are widely utilized in traditional recommendation model, they are naturally treated as an item-indexing way for aligning LLMs with recommendation tasks. For description-based identifiers, researchers generally use the title or other textual attributes of the item as an identifier, and formalize it into natural language as input (Dai et al., 2023a; Liu et al., 2023c). Due to the richness of natural language, description-based identifiers have high readability and flexibility. For hybrid identifiers, considering collaborative and textual information of users or items, LLMs can perceive both token-based and description-based identifiers (Liao et al., 2023; Li et al., 2023b). Current work using hybrid identifiers can be divided into different types. Some of them add item sequence with textual description and ID information respectively into different positions of prompts (Zhang et al., 2023b), while others combine token-based and description-based embeddings together for each item in sequences through concatenation (Li et al., 2023b).
@@
$\bullet$ Grounding to Candidate Items. After constructing candidate items for recommendation, how to ground the output of LLMs into recommendation item lists is a crucial problem. Methods for grounding to candidate items are associated with different types of recommendation tasks. For discriminative recommendation, which provide candidate item lists in the prompts, it is easy for LLMs to follow instructions and generate the output right within the candidate item lists through elaborately designed prompts (Zhiyuli et al., 2023; Liao et al., 2023). For generative recommendation tasks, which do not explicitly provide candidate item list in prompts but require further processing for the output, recent work indicates effective approaches for mapping the output of LLMs into candidate item lists (Zheng et al., 2023). For example, GPT4Rec (Li et al., 2023m) first generates hypothetical “search queries” using titles of recently interacted items, and then retrieves similar items from the item pool by BM25 algorithm. E4SRec (Li et al., 2023b) directly utilizes the nearest neighbor search between the output of LLMs and the vectors in the item linear projection component for grounding to candidate items. In this case, the recommendation process will not suffer from length limitation problems since they choose not to include candidate item lists into the input of LLMs.
@@
As shown in Figure 1, our framework ProLLM4Rec involves two mapping processes between the recommendation space and language space: (1) when using LLMs to represent users, items in the recommendation space need to be used as historical sequences. (2) When re-ranking candidate items, it is required to provide candidate items in the recommendation space for LLMs. The key issue of semantic space alignment is the indexing and grounding of items. As for the item indexing, pseudo ID-based sequences and description-based natural languages are two typical forms. As for the grounding methods, the generative method renders LLMs to output labels of all candidate items, and other mapping strategies such as logits distribution and similarity calculation can also be considered (Yue et al., 2023; Li et al., 2023b).
@@ 5.3.2. Research Questions and Experimental Setup
In this section, we explore the effect of candidate items on ProLLM4Rec, and provide empirical findings on the following research questions:
  • RQ1: Whether different representations of candidate items lead to similar recommendation performance?
@@
In this paper, we explore the performance differences between the following two item representation methods:
  • ID-based identifiers: we can assign serial numbers or indicators (e.g., ABCD) to the candidate items, and instruct LLMs to only output the re-ranked indicators instead of the complete text.
  • Description-based identifiers: we instruct LLMs to directly output the complete text e.g., titles of candidate items, and the re-ranking results are parsed from the textual output from LLMs.
Moreover, researchers tend to adopt a two-stage recommendation paradigm, in which conventional recommenders initially recall multiple candidates, and subsequently, LLMs evaluate and rank these candidates to provide enhanced recommendations. This approach effectively reduces the computational load of LLMs, enabling more efficient deployment. However, there is still no consensus regarding the effectiveness of LLMs in reranking the candidates retrieved by traditional recommendation models (Hou et al., 2023; Ren et al., 2023). In order to address this, we conduct experiments to compare the impact of LLMs in re-ranking the original recommendation results generated by traditional recommenders. We evaluate three classical recommenders, including Pop, BPR (Rendle et al., 2009), and SASRec (Kang and McAuley, 2018), and report the result in Table 11.
Table 10. Performance comparison on the grounding forms of LLMs w.r.t. candidate items.
Method MovieLens-1M Amazon-Books
ndcg@1 ndcg@10 ndcg@20 ndcg@1 ndcg@10 ndcg@20
ID-based identifiers 0.2750 0.5423 0.5725 0.0600 0.1473 0.1685
@@
Table 11. Performance comparison on the effects of LLMs w.r.t. the source of candidate items.
Method MovieLens-1M Amazon-Books
ndcg@1 ndcg@10 ndcg@20 ndcg@1 ndcg@10 ndcg@20
@@ 5.3.3. Observations and Discussion
As shown in Table 10-11, findings of candidate items construction are as follows.
(1) Grounding forms of candidate items (RQ1). Firstly, we focus on the grounding forms of candidate items. As shown in Table 10, we can see that whether in the movie dataset or the book dataset, the complete name of the output item is more convenient for extracting and grounding recommendation results than the specified identifier. In other words, description-based identifiers outperform ID-based identifiers when grounding candidate items for LLMs (Zhao et al., 2023; Hou et al., 2023). However, whether it is ID-based or description-based, the inference time of the grounding strategy based on generative output will increase with the increasing number of candidate items. A feasible solution is to use one prediction to obtain the probability score on all candidate items based on the output logits (Yue et al., 2023). Language models are more sensitive to the textual output rather than pure identifiers, but we can also use more advanced index strategies and grounding methods to improve the accuracy of text and item mapping (Hua et al., 2023), which requires continuous exploration by future research (Ren et al., 2023).
(2) The evaluation for re-ranking abilities of LLMs (RQ2). Secondly, we discuss the performance difference between traditional algorithms and re-ranking results after LLMs. As shown in Table 11, LLMs improve results of the traditional recommendation method BPR on two datasets by a large margin. However, BPR is not specifically designed for sequential recommendation and its original performance is not good. As for the basic method Pop and the typical sequential recommendation model SASRec, the re-ranking effect of LLMs varies between the two datasets. For the MovieLens-1M dataset in the movie domain, our initial prompts cannot instruct LLMs to achieve better results than traditional methods such as Pop and SASRec. The possible reason is that this movie dataset is very dense, and fully-trained collaborative filtering signals are more important than general knowledge of movies. While in the Amazon-Books dataset, the re-ranking of LLMs can further improve results of Pop and SASRec, possibly because the general knowledge of LLMs for books can effectively adapt to the sparse recommendation data. The overall experimental results indicate that the re-ranking performance of LLMs for recommendation results varies depending on specific recommendation methods and datasets. However, it is acknowledged that the general knowledge of LLMs can complement the collaborative signals of traditional models, tapping the potential to improve traditional recommendation algorithms (Li et al., 2023a; Zhang et al., 2023b).
Observations of Candidate Items Construction Candidate items construction for LLMs is crucial to the final recommendation results. Retrieving candidate items by traditional recommendation models first, and then re-ranking items by LLMs can further improve the results, but the performance varies on specific methods and datasets.

### v2

@@ 5.2.5. Observations and Discussion
As shown in Table 9 and Table 10, we summarize the following findings for modeling the user interest in LLM-RS:
@@
(1) The design of query and memory for retrieval-based modeling methods (RQ1). We conduct an experiment to explore the retrieval strategies for recent and relevant items, and evaluate the selection of query and memory for item retrieval. As shown in Table 10, we compare four combinations with queries and memories. In terms of the query, combining the design of long-term and short-term interest yields better results than only the short-term interest. When selecting items that represent interest for recommendation, the query needs to consider both recent tendencies and long-term preferences of users. Our results have shown that the personalized long-term interest of users can have a positive effect on retrieval for interest modeling. In terms of the memory, the design of the personalized memory with user interest outperforms the global memory with general descriptions. This result is in line with features of personalized recommender systems (Shu et al., 2023; Wang et al., 2024c), indicating the importance of personalization when designing the user interest for LLMs (Zhang et al., 2024a).
(2) The impact of different modeling forms of the user interest (RQ2). In Table 9, we compare the user interest modeling forms across three modeling methods (memory-based, retrieval-based and generation-based), and highlight three interest types (short-term, long-term and hybrid interest) with three colors. From the perspective of modeling methods, the recommendation performance of retrieval-based and generation-based methods is generally better than that of memory-based methods, indicating the importance of Retrieval-Augmented Generation (RAG) from the interest memory. Meanwhile, on the MovieLens-1M dataset with dense interactions, generative summarization based on short-term interest can further improve the effect, while on the Amazon books dataset with sparse interactions, retrieval-based modeling methods achieve the best, indicating that the retrieval results cannot be completely replaced by summative sentences. On the other hand, long-term interest and short-term interest can complement each other, and combining the two may yield better results. Instead of concatenating the long-term and short-term sequences, one possible combination approach is to use the short-term interest as queries to retrieve the long-term interest, serving as the domain knowledge for LLMs to model user preferences.
Observations of Modeling User Interest While selecting recent and relevant items, it is preferable to combine long-term and short-term user profiles with personalized descriptions for query and memory. The recommendation performance of retrieval-based and generation-based methods is generally better than that of memory-based methods, indicating the importance of Retrieval-Augmented Generation (RAG) from the interest memory. Long-term interest and short-term interest can complement each other, and combining the two may yield better results. Retrieving long-term preferences based on short-term intentions is a suitable way for forming hybrid interest descriptions, and more approaches can be further explored.
@@ 5.3.1. Procedures of Candidate Items Construction
When leveraging LLMs as recommender systems, a crucial problem is that LLMs tend to recommend items not in the dataset (Geng et al., 2022). One of the approaches is to provide a limited number of candidate items for model selection. On the one hand, the source of candidate items determines the upper limit of personalized recommendation accuracy (Hou et al., 2024; Zhang et al., 2023e). On the other hand, the form of candidate items reflects the specific recommendation task and exerts huge impacts on performance of LLMs for recommendation (Dai et al., 2023a). As shown in Fig. 10, we will discuss the main procedures of processing candidate items, including the source, representation and grounding of candidate items, and then conduct experiments on the effects of candidate items.
@@
Figure 10. Main procedures of candidate items construction in  LLM-RS, including the source, representation and grounding of candidate items.
[…]
Main procedures of candidate items construction in  LLM-RS, including the source, representation and grounding of candidate items.
@@
$\bullet$ Source of candidate items. In practical industrial applications, recommendation is a multi-stage process, including two kinds of typical stages, i.e., the recall stage and the re-ranking stage (Bao et al., 2023a; Guo et al., 2017; Rendle et al., 2009). The recall stage in recommender systems is used for preliminary filtering and selection from the entire set of item candidates, while the re-ranking stage is used for readjusting the selected candidates, which is more suitable for LLMs as the re-ranker (Hou et al., 2024; Ma et al., 2023). Therefore, the source of candidate items to be ranked is crucial for the results of LLMs for recommendation. Popular approaches for selection of candidate items include traditional recommendation models (Yue et al., 2023) and retrieval algorithms (Lin et al., 2024a), both of which provide a list of potentially recommended items.
@@
$\bullet$ Representation of candidate items. In the construction of prompts, we need to use historical items to represent users and provide candidate items for LLMs to rank, both of which involve item indexing (Geng et al., 2022; Lin et al., 2023b). However, there is a gap between the item representation in LLMs and recommendation, so the indexing of items between LLMs and recommender systems becomes the key and difficult point of LLM-RS. There are three typical ways to index items: 1) token-based identifiers, 2) description-based identifiers and 3) hybrid identifiers (Zheng et al., 2024; Wang et al., 2024a). For token-based identifiers, researchers often use numerical IDs to identify items (Geng et al., 2022). Since token-based identifiers are widely utilized in traditional recommendation model, they are naturally treated as an item-indexing way for aligning LLMs with recommendation tasks. For description-based identifiers, researchers generally use the title or other textual attributes of the item as an identifier, and formalize it into natural language as input (Dai et al., 2023a; Liu et al., 2023c; Bao et al., 2023a). Due to the richness of natural language, description-based identifiers have high readability and flexibility. For hybrid identifiers, considering collaborative and textual information of users or items, LLMs can perceive both token-based and description-based identifiers (Liao et al., 2023; Li et al., 2023c). Current work using hybrid identifiers can be divided into different types. Some of them add item sequence with textual description and ID information respectively into different positions of prompts (Zhang et al., 2023b), while others combine token-based and description-based embeddings together for each item in sequences through concatenation (Li et al., 2023c).
@@
$\bullet$ Grounding to candidate items. After constructing candidate items for recommendation, how to ground the output of LLMs into recommendation item lists is a crucial problem. Methods for grounding to candidate items are associated with different types of recommendation tasks. For discriminative recommendation, which provide candidate item lists in the prompts, it is easy for LLMs to follow instructions and generate the output right within the candidate item lists through elaborately designed prompts (Li et al., 2023b; Liao et al., 2023). For generative recommendation tasks, which do not explicitly provide candidate item list in prompts but require further processing for the output, recent work indicates effective approaches for mapping the output of LLMs into candidate item lists (Zheng et al., 2024). For example, GPT4Rec (Li et al., 2023k) first generates hypothetical “search queries” using titles of recently interacted items, and then retrieves similar items from the item pool by BM25 algorithm. E4SRec (Li et al., 2023c) directly utilizes the nearest neighbor search between the output of LLMs and the vectors in the item linear projection component for grounding to candidate items. In this case, the recommendation process will not suffer from length limitation problems since they choose not to include candidate item lists into the input of LLMs.
@@
As shown in Figure 1, our framework involves two mapping processes between the recommendation and language space: (1) when using LLMs to represent users, items in the recommendation space need to be used as historical sequences. (2) When re-ranking candidate items, it is required to provide candidate items in the recommendation space for LLMs. The key issue of semantic space alignment is the indexing and grounding of items. As for the item indexing, token-based sequences and description-based natural languages are two typical forms. As for the grounding methods, the generative method renders LLMs to output labels of all candidate items, and other mapping strategies such as logits distribution and similarity calculation can also be considered (Yue et al., 2023; Li et al., 2023c).
@@ 5.3.2. Research Questions and Experimental Setup
In this section, we explore the effect of candidate items on LLM-RS, and provide empirical findings on the following research questions:
  • RQ1: Whether representations of items lead to similar recommendation performance?
@@
We explore the performance differences between the following two item representation methods. Basic prompts used in this section is illustrated in Fig. 2(a), and ChatGPT is employed as the zero-shot re-ranker.
  • Token-based identifiers: we can assign serial numbers or indicators (e.g., ABCD) to the candidate items, and instruct LLMs to only output the re-ranked identifiers.
  • Description-based identifiers: we instruct LLMs to directly output the complete text e.g., titles of candidate items, and the re-ranking results are parsed from the textual output.
Moreover, researchers tend to adopt a two-stage recommendation paradigm, in which conventional recommenders initially recall multiple candidates, and subsequently, LLMs evaluate and rank these candidates to provide enhanced recommendations. This approach effectively reduces the computational load of LLMs, enabling more efficient deployment. However, there is still no consensus regarding the effectiveness of LLMs in reranking the candidates retrieved by traditional recommendation models (Hou et al., 2024; Ren et al., 2024). In order to address this, we conduct experiments to compare the impact of ChatGPT in re-ranking the original recommendation results generated by traditional recommenders. We evaluate three classical recommenders, including Pop, BPR (Rendle et al., 2009), and SASRec (Kang and McAuley, 2018).
Table 11. Performance comparison on the grounding forms of LLMs w.r.t. candidate items. Token-based identifiers represent items numerically, while the description-based identifiers denote textual titles.
Method MovieLens-1M Amazon Books
N@1 N@10 N@20 N@1 N@10 N@20
Token-based identifiers 0.2750 0.5423 0.5725 0.0600 0.1473 0.1685
@@
Table 12. Performance comparison of traditional recommendation methods with and without ChatGPT as a re-ranker. “+ LLM” denotes the recommendation results of LLMs by re-ranking the top-20 candidate items of traditional models. “Impr.” denotes the percentage improvement (or degradation) achieved by LLMs.
Method MovieLens-1M Amazon Books
N@1 N@10 N@20 N@1 N@10 N@20
@@ 5.3.3. Observations and Discussion
Findings of candidate items construction are as follows.
(1) Grounding forms of candidate items (RQ1). Firstly, we focus on the grounding forms of candidate items. As shown in Table 11, we can see that whether in the movie dataset or the book dataset, the complete name of the output item is more convenient for extracting and grounding recommendation results than the specified identifier. In other words, description-based identifiers outperform Token-based identifiers when grounding candidate items for LLMs (Zhao et al., 2023; Hou et al., 2024). However, whether it is ID-based or description-based, the inference time of the grounding strategy based on generative output will increase with the increasing number of candidate items. A feasible solution is to use one prediction to obtain the probability score on all candidate items based on the output logits (Yue et al., 2023). Language models are more sensitive to the textual output rather than pure identifiers, but we can also use more advanced index strategies and grounding methods to improve the accuracy.
(2) The evaluation for re-ranking abilities of LLMs (RQ2). Secondly, we discuss the performance difference between traditional algorithms and re-ranking results after LLMs. As shown in Table 12, LLMs improve results of the traditional recommendation method BPR on two datasets by a large margin. As for the basic method Pop and SASRec, the re-ranking effect of LLMs varies between the two datasets. For the MovieLens-1M dataset in the movie domain, our initial prompts cannot instruct LLMs to achieve better results than traditional methods such as Pop and SASRec. The possible reason is that this movie dataset is very dense, and fully-trained collaborative filtering signals are more important than general knowledge of movies. While in the Amazon Books dataset, the re-ranking of LLMs can further improve results of Pop and SASRec, possibly because the general knowledge of LLMs for books can effectively adapt to the sparse recommendation data. The overall experimental results indicate that the re-ranking performance of LLMs for recommendation results varies depending on specific recommendation methods and datasets. However, it is acknowledged that the general knowledge of LLMs can complement the collaborative signals of traditional models, tapping the potential to improve traditional recommendation algorithms (Li et al., 2023a; Zhang et al., 2023b).
Observations of Candidate Items Construction LLMs may exhibit limitations in retaining critical information regarding the candidate list from prompt instructions, potentially leading to responses that deviate from the intended requirements. Description-based identifiers outperform token-based identifiers when grounding candidate items. Retrieving candidate items by traditional recommendation models first, and then re-ranking items by LLMs can further improve the results.